To what volume in ml would you have to change 55.0 ml of gas at 201.4 kpa in order to decrease its pressure to 17.0 kpa is temperature and number of molecules remain constant
1
Expert's answer
2019-04-08T07:55:06-0400

P × V = n × R × T


T = const, n = const.

Then P1 × V1 = P2 × V2

201.4 kPa × 55.0 mL = 17.0 kPa × V2

V2 = 201.4 kPa × 55.0 mL / 17.0 kPa = 651.59 mL ≈ 651.6 mL.